OpenStudy (anonymous):

What would I do to solve this equation -x + 3y=6

OpenStudy (mchilds15):

Well you have to isolate the So what would be the first step you take?

OpenStudy (mchilds15):

**isolate the x

OpenStudy (mathmale):

Depends upon what you mean by "solve." Seriously: you have two unknowns here (x and y). With only one equation, the best you can do would be to solve for either x or y in therms of the other.

OpenStudy (anonymous):

In the question it says to find the x-intercept ?

OpenStudy (mchilds15):

Ahhhh well then this is a breeze hun :)

OpenStudy (anonymous):

I really suck at algebra and I've been out of school a lot

OpenStudy (mchilds15):

Hey mathmale, how would we do this again? Isn't there a certain name for the way you work this problem?

OpenStudy (texaschic101):

to find the x intercept of an equation, sub in 0 for y and solve for x. -x + 3y = 6 -- sub in 0 for y -x + 3(0) = 6 -x = 6 -- multiply by -1 to make x positive x = -6 your x intercept is (-6,0)

OpenStudy (texaschic101):

if you were looking for the y intercept, you could just sub in 0 for x and solve for y

OpenStudy (anonymous):

Thank you so much😃

OpenStudy (texaschic101):

sure thing :)